OpenStudy (anonymous):

Charlene has nine different necklaces, and she wears four at a time. How many different ways can Charlene wear four necklaces?

OpenStudy (anonymous):

\[\dbinom{9}{4}\]

OpenStudy (anonymous):

compute via \[\frac{9\times 8\times 7\times 6}{4\times 3\times 2}\] \[=9\times 2\times 7=216\]

OpenStudy (anonymous):

or if you have a nice calculator use \[9C4\]

OpenStudy (anonymous):

thanks again :)